# Dual Modalities of Text: Visual and Textual Generative Pre-training

### v2

@@ Abstract
Harnessing visual texts represents a burgeoning frontier in the evolution of language modeling. In this paper, we introduce a novel pre-training framework for a suite of pixel-based autoregressive language models, pre-training on a corpus of over 400 million documents rendered as RGB images. Our approach is characterized by a dual-modality training regimen, engaging both visual data through next patch prediction with a regression head and textual data via next token prediction with a classification head. This study is particularly focused on investigating the synergistic interplay between visual and textual modalities of language. Our comprehensive evaluation across a diverse array of benchmarks reveals that the confluence of visual and textual data substantially augments the efficacy of pixel-based language models. Notably, our findings show that a unidirectional pixel-based model, devoid of textual data during training, can match the performance levels of advanced bidirectional pixel-based models on various language understanding benchmarks. This work highlights the considerable untapped potential of integrating visual and textual information for language modeling purposes. We will release our code, data, and checkpoints to inspire further research advancement.
Dual Modalities of Text: Visual and Textual Generative Pre-training
Yekun Chai  Qingyi Liu∗♡  Jingwu Xiaothanks: Work done during QL and JX’s internship at Baidu. Shuohuan Wang   Yu Sun   Hua Wu Baidu Inc.   Sun Yat-sen University   Peking University {chaiyekun,wangshuohuan,sunyu02}@baidu.com {liuqy95}@mail2.sysu.edu.cn
@@ 1 Introduction
The landscape of large language models (LLMs) is undergoing a significant transformation, with advancements that extend the boundaries of language assistant Touvron et al. (2023a), code generation Lozhkov et al. (2024); Chai et al. (2023), and multimodal comprehension OpenAI (2023); Anil et al. (2023). These models traditionally tokenize input data into discrete elements, treating them as sequences of identifiers, thereby enabling diverse applications. However, this approach often struggles with visually enriched textual content, such as PDFs, where direct parsing into text incurs significant information loss. Traditional methodologies typically employ pre-trained optical character recognition (OCR) tools for extracting information from such visual texts, but these methods are inherently limited by the fidelity of text extraction.
In response to these challenges, a novel paradigm of pixel-based language modeling has emerged, offering a direct pathway to learning from text as visual data (images), transcending the constraints of textual modality Rust et al. (2023); Tschannen et al. (2023). This approach promises to surmount the vocabulary bottleneck issue Rust et al. (2023)—a trade-off inherent in balancing input encoding granularity against the computational feasibility of vocabulary probability estimation in conventional language models.
@@
Moreover, current research often processes visual text as 8-bit grayscale Rust et al. (2023) or 2-bit binary images Tai et al. (2024). This approach restricts the representation of color, critical for elements like emojis and font highlights, and diverges from the natural image format in RGB. Notably, there appears to be a lack of studies pre-training on RGB images, which could more accurately reflect the complexities of visual text.
This research aims to fill these gaps by offering a comprehensive examination of the effects of pixel-based versus text-based pre-training within an autoregressive language modeling context. Our study is steered by three critical research questions:
RQ1: Feasibility of pure pixel-based autogressive pre-training on RGB images of visual texts. Can an autoregressive language model trained solely on 24-bit RGB images of visual texts achieve competitive performance?
RQ2: Impact of autoregressive pixel pre-training on multilingual tasks. We explore whether autoregressive pixel pre-training can overcome the vocabulary bottleneck in multilingual contexts, assessing its effectiveness in generalizing linguistic features across languages.
RQ3: Synergistic effects of multimodal pre-training. How do pixel-based and text-based pre-training synergize, and in what ways does this multimodal strategy enhance the model’s performance on language understanding tasks and its cross-lingual applicability?
Contributions   (1) We introduce a novel autoregressive pre-training approach that combines the pre-training objective of next token and next patch prediction, bridging text and visual modalities through tailored classification and regression heads, respectively. This methodology marks a significant advancement in multimodal language model training. (2) We construct a comprehensive visual text dataset of over 400 million documents by text rendering for pixel-based pre-training (equivalent to roughly 236 billion text tokens). We will release the fine-tuning datasets used for language understanding and multilingual evaluation in our experiments, facilitating further advancements in this emerging field. (3) In this work, we demonstrate that pre-training decoder-only transformers on RGB visual images can surpass or match the performance of encoder-based models in language understanding and multilingual transfer tasks. Our findings suggest that leveraging the rich information content of RGB images in decoder-only architectures not only enhances model understanding of complex linguistic patterns but also facilitates effective knowledge transfer across languages. (4) We empirically demonstrate the substantial potential of integrating visual and textual data for enhanced language model training.
@@ 2 Related Work
Pixel Representations for Text Advances in pixel-based language modeling have increasingly focused on exploiting the orthographic and typographic properties of text through visual representations. PIXEL Rust et al. (2023) utilizes masked auto-encoders to address the vocabulary bottleneck by reconstructing pixels in masked text images. Moreover, CLIPPO Tschannen et al. (2023) demonstrates enhanced language comprehension using a unified encoder for both image and text modalities. Further research by Lotz et al. (2023) evaluates the impact of rendering techniques on the efficacy of pixel-based encoders. These studies primarily utilize bidirectional encoders and process text as grayscale images.
@@
Autoregressive Pre-training on Pixels   Existing methods in pixel-based autoregressive pre-training divide into vector quantization techniques—transforming continuous images into discrete tokens—and direct pixel prediction. These approaches include VQ-VAE Van Den Oord et al. (2017) and VQGAN Esser et al. (2021) followed by next token prediction Chen et al. (2020); Ramesh et al. (2021), and prefix language modeling that predicts future visual patches from bidirectional pixel contexts El-Nouby et al. (2024).
@@ 3.2 Input Representation
Image Input Inspired by the Vision Transformer (ViT, Dosovitskiy et al., 2020), our method tailors the image patch processing paradigm to the sequential processing needs of autoregressive transformer decoders handling visual text imagery, as shown in Figure LABEL:fig:pixelgpt. This process commences by rendering the textual input as RGB images $x\in\mathbb{R}^{H\times W\times C}$ as aforementioned in §3.1, subsequently partitioning these into uniform patches $x_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$ illustrated as Figure 8, where $(H,W)$ defines the original image’s resolution, $(P,P)$ specifies each patch’s resolution with $P=H$, and $N=W/P$ denotes the total number of patches. The patches are then flattened, mapped to a $D$-dimensional space through a learnable linear projection, and finally fed into the transformer’s sequential processing stream. Unlike ViT, which caters to two-dimensional inputs, our model processes these patches in the sequence order in which the text appears, emulating the linear progression of reading. This patch-based segmentation aligns with the sequential nature of language, enabling our model to predictively learn from the visual data.
Text Input We leverage the same tokenizer as Llama 2, segmenting input text into discrete tokens with a total vocabulary size of 32k. These tokens are then transformed into dense vector representations through an embedding lookup table.
@@ 3.3 Pre-training Objectives
Next Patch Prediction  Given a sequence of $N$ visual patches $x_{p}=(x^{1}_{p},x^{2}_{p},\cdots,x^{N}_{p})$ where each visual patch $x^{t}_{p}$ is a flattened patch embedding. We decompose the image patch sequence into the production of $N$ conditional probabilities:
$p(x^{1}_{p},x^{2}_{p},\cdots,x^{N}_{p})=\prod_{t=1}^{N}p(x^{t}_{p}|x^{1}_{p},x^{2}_{p},\cdots,x^{t-1}_{p})$ (1)
@@
Next Token Prediction  For text inputs, we utilize a conventional next token prediction objective, optimizing a cross-entropy loss that evaluates the fidelity of predicted token sequences generated via teacher-forcing against the ground truth tokens.
@@ 3.5 Pre-training Details
Model Architecture Our architecture, illustrated in Figure LABEL:fig:arch, is built upon a stack of $N=24$ standard transformer decoder Vaswani et al. (2017), following Llama 2 Touvron et al. (2023b). We incorporate RMSNorm for pre-normalization Zhang and Sennrich (2019), SwiGLU activation functions Shazeer (2020); Chai et al. (2020), rotary position embeddings Su et al. (2024), and grouped query attention Ainslie et al. (2023). Comprehensive specifications and additional implementation details of our architecture are in Appendix §B.
Data For visual image data, we use rendered the corpus of peS2o, English Wikipedia and C4 datasets for pre-training; while for text data, we adopt peS2o, English Wikipedia, C4, Common Crawl, and The Stack v1. We refer the readers to Appendix §C for details.
@@ 4.1 Experimental Setup
Baselines  For a thorough evaluation, we benchmark against models specialized in textual and visual representations. In the textual category, BERT and GPT-2 Radford et al. (2019) are chosen. For pixel-based models, we contrast our approach with DONUT Kim et al. (2022), CLIPPO Tschannen et al. (2023), PIXEL Rust et al. (2023), and PIXAR Tai et al. (2024), which are trained on pixel-based representation. Detailed discussions are provided in Appendix §F.
@@ 4.2 Results
RQ1: Autoregressive Pixel-based Pre-training Rivals PIXEL.
Our empirical investigation, detailed in Table 1, scrutinizes the feasibility of pure pixel-based autoregressive pre-training on RGB images of visual texts. The proposed PixelGPT model, training solely on rich 24-bit RGB visual inputs, demonstrates not merely a competitive edge but, in several tasks, surpasses the performance of models pre-trained on text alone. Specifically, PixelGPT exhibits remarkable superiority on GLUE benchmarks—evidenced by its marked performance increases on the STS-B (+5.4), MRPC (+13.1), RTE (+11.9), and WNLI (+4.3) assessments compared to GPT-2. This demonstrates the viability of pixel-based pre-training in capturing complex linguistic constructs.
Moreover, when juxtaposed with PIXEL, which leverages a bidirectional encoder architecture, PixelGPT exhibits enhanced performance in QQP (+1.5), RTE (+3.4), and WNLI (+5.4). These results collectively affirm the hypothesis that autoregressive pre-training on RGB visual data is feasible and advantageous for language modeling. PixelGPT achieves the optimal performance among pixel-based approaches on GLUE, underscoring the transformative impact of integrating rich visual information into pre-training.
@@
RQ2: Impact of Autoregressive Pixel Pre-training on Multilingual Tasks.
RQ3: Synergistic Effects of Multimodal Pre-training.
@@
This comparative analysis, spanning both GLUE and XNLI datasets (the latter within the translate-train-all settings), is shown in Tables 3 and 4. A pivotal finding is that the incorporation of dual-modality data during pre-training markedly enhances average performance across language understanding tasks: DualGPT (76.9) surpasses both TextGPT (76.3) and MonoGPT (75.4). This suggests that potential conflicts arising from unimodal training can be significantly alleviated through a multimodal pre-training approach. This inference is corroborated by XNLI outcomes, wherein the addition of pixel-text paired data bolstered the model’s multilingual interpretative proficiency.
@@ 4.3 Analysis
Scaling Training Tokens vs. GLUE Performance  In Figure 3, we delineate the correlation between the scale of training data and the ensuing performance on the GLUE benchmark. Our analysis encompasses a spectrum of total training tokens/patches from 10 billion (B) to 240B, juxtaposing the trajectories of TextGPT, PixelGPT, MonoGPT, and DualGPT, with BERT and PIXEL serving as benchmarks. The MonoGPT and DualGPT models are evaluated under two different input modalities: text and pixel. From our findings, two primary insights emerge: (1) Pixel-based autoregressive pretraining models exhibit an increased data demand. With minimal training (e.g., at 10B), pixel-based models initiate at a lower performance threshold in pixel modality (all under 55%), compared to their text modality counterparts, which approximate a performance level of 70%. Nevertheless, with the increase of training data, a critical volume threshold catalyzes a substantial rise in performance for PixelGPT, MonoGPT, and DualGPT in pixel modality. This trajectory reveals a progressive convergence of PixelGPT towards the text-based baseline, culminating in its overtaking of PIXEL at around 200B tokens/patches and nearing TextGPT with a less than 5-point performance differential, while still on an upward trend. (2) The integration of paired dual-modality data during pretraining appears to confer significant benefits on multimodal learning, particularly for pixel-based input. When matched for training data volume, DualGPT consistently eclipses MonoGPT across comparable benchmarks, with the former maintaining a pronounced lead in pixel modality. This trend underscores the value of incorporating paired text-image data in pretraining to enhance the efficacy of multimodal learning.
@@
Scaling Training Tokens vs. XNLI (Translate-Train-All) Performance We further explored the progression of model performance in multilingual capability across varying volumes of pre-trained tokens/patches. This comparison, delineated in Figure 4, focused on the Translate-Train-All setting of the XNLI benchmark. (1) Pixel-based autoregressive models display a heightened requirement for training data in multilingual tasks, corroborating the trend observed on the GLUE benchmark. Initially, there is a notable performance disparity between pixel and text modalities, with pixel-based models lagging behind when training on a lesser volume of tokens/patches. However, this gap diminishes substantially with the increase in training volume. Remarkably, upon reaching the 200B, PixelGPT not only surpasses PIXEL but also matches the performance of BERT, indicating a continued potential for further enhancement in its multilingual proficiency with additional training data. (2) The injection of dual-modality data at the early stages of training appears to be particularly beneficial for models learning from pixel data. When comparing DualGPT and MonoGPT under the pixel modality, DualGPT demonstrates a notable performance advantage at the outset of training (55% vs. 45.8% at the 10B token/patch mark). Although this edge tapers as the training volume expands, it suggests that early-stage multimodal alignment aids the pixel-based models in leveraging the textual data for enhanced multilingual understanding. (3) Our text-based pre-training approach, TextGPT, demonstrates superior results over BERT. This is evident when training reaches approximately 100B tokens, where TextGPT outperforms BERT. This improvement may be attributed, in part, to our byte-level BPE tokenization as utilized in Llama 2, which effectively deconstructs unseen languages into their constituent raw bytes—a capability not afforded by BERT. Additionally, the enrichment of our text pre-training corpus from diverse sources contributes to this. For a detailed breakdown of the text pre-training data, we refer readers to Appendix §C.2.
A Large Batch Size Improves Stable Training  We observe a distinct preference for larger batch sizes when fine-tuning pixel-based modalities across certain datasets. As in Figure 5, we evaluate how different batch sizes—64, 128, 256, and 512—affect model performance on selected GLUE benchmark tasks, namely QQP, CoLA, and STS-B. A clear trend emerges from the data: increasing the batch size correlates with improved model performance. Our analysis suggests that pixel modality fine-tuning exhibits greater variance than text modality and benefits from the use of larger batch sizes. This appears to mitigate the variability inherent in different training batches, thus enhancing training stability. It prevents premature convergence to suboptimal local minima and fosters higher model accuracy.

### v3

# Autoregressive Pre-Training on Pixels and Texts

@@ Abstract
The integration of visual and textual information represents a promising direction in the advancement of language models. In this paper, we explore the dual modality of language—both visual and textual—within an autoregressive framework, pre-trained on both document images and texts. Our method employs a multimodal training strategy, utilizing visual data through next patch prediction with a regression head and/or textual data through next token prediction with a classification head. We focus on understanding the interaction between these two modalities and their combined impact on model performance. Our extensive evaluation across a wide range of benchmarks shows that incorporating both visual and textual data significantly improves the performance of pixel-based language models. Remarkably, we find that a unidirectional pixel-based model trained solely on visual data can achieve comparable results to state-of-the-art bidirectional models on several language understanding tasks. This work uncovers the untapped potential of integrating visual and textual modalities for more effective language modeling. We release our code, data, and model checkpoints at https://github.com/ernie-research/pixelgpt.
Autoregressive Pre-Training on Pixels and Texts
Yekun Chai  Qingyi Liu∗♡  Jingwu Xiaothanks: Work done during QL and JX’s internship at Baidu. Shuohuan Wang   Yu Sun   Hua Wu Baidu Inc.   Sun Yat-sen University   Peking University {chaiyekun,wangshuohuan}@baidu.com {liuqy95}@mail2.sysu.edu.cn
@@ 1 Introduction
Recent advancements in large language models (LLMs) have pushed the boundaries of their capabilities in diverse applications, including language assistant Touvron et al. (2023a), code generation Lozhkov et al. (2024); Chai et al. (2023), and multimodal comprehension OpenAI (2023); Anil et al. (2023). LLMs typically tokenize input text into sequences of discrete subword units, allowing for a wide array of applications. However, tokenization-based approaches struggle with visually complex textual content, such as PDFs, where converting visual data into plain text often results in significant information loss. Traditional solutions rely on optical character recognition (OCR) models for extracting text from images, but these methods are inherently limited by the accuracy of text extraction and the fidelity of the original document structure.
To address these challenges, recent work has introduced a new paradigm: pixel-based language modeling. This approach learns directly from the visual representation of text (as images) rather than relying solely on tokenized text. Models such as PIXEL Rust et al. (2023) exemplify this shift, offering solutions that circumvent the limitations of traditional tokenization by treating text as image data. Pixel-based modeling also addresses the vocabulary bottleneck—a trade-off between input encoding granularity and the computational costs associated with vocabulary estimation in conventional language models Rust et al. (2023).
@@
Moreover, current research often processes visual text as 8-bit grayscale Rust et al. (2023) or 2-bit binary images Tai et al. (2024). This approach constrain the richness of the visual input, especially when processing content with color information, such as emojis or highlighted text. This limitation suggests that processing real-valued RGB images could offer a more detailed representation of visual text. However, the potential of pre-training autoregressive language models on raw RGB images, which more closely mirror the natural visual characteristics of documents, has not been fully explored.
This research addresses two distinct challenges in language modeling: (1) the feasibility of tokenization-free autoregressive pre-training using PixelGPT, and (2) the synergistic benefits of multimodal pre-training with DualGPT.
First, we focus on the performance of PixelGPT, a tokenization-free model that processes raw visual text images. We investigate whether training an autoregressive model directly on real-valued pixels can achieve competitive results without tokenization, particularly in multilingual contexts. This exploration assesses whether PixelGPT can overcome the vocabulary bottleneck in multilingual tasks by generalizing linguistic features across diverse languages, thus bypassing the constraints of predefined vocabularies typically encountered in traditional text-based models.
Second, we evaluate DualGPT, which integrates both visual and textual modalities during pre-training. By leveraging pixel-based and text-based pre-training together, DualGPT is designed to harness the interaction between these two modalities. We explore how this multimodal strategy improves model performance on language understanding tasks and cross-lingual generalization, offering advantages over models that rely on a single modality.
Contribution
To conclude, our main contributions are as follows:
  • We empirically demonstrate the substantial potential of integrating visual text images for enhanced language model training. We show that pre-training decoder-only transformers on visual images can match or slightly underperform compared to text-based inputs but achieve competitive results with bidirectional PIXEL models Rust et al. (2023). This illustrates the potential for scaling trends to eventually surpass text-based pre-trained models.
  • We systematically explore autoregressive pre-training on both visual text images and plain text modalities, demonstrating the potential of causal language models to effectively learn from visual text images and highlighting the interplay between different modalities.
  • We release our fine-tuning datasets for language understanding and multilingual evaluation [^1] https://github.com/ernie-research/pixelgpt, facilitating further research in this emerging field.
@@ 2 Related Work
2.1 Pixel Representations for Text
Advances in pixel-based language modeling have increasingly focused on exploiting the orthographic and typographic properties of text through visual representations. PIXEL Rust et al. (2023) utilizes masked auto-encoders to address the vocabulary bottleneck by reconstructing pixels in masked text images. Moreover, CLIPPO Tschannen et al. (2023) demonstrates enhanced language comprehension using a unified encoder for both image and text modalities. Further research by Lotz et al. (2023) evaluates the impact of rendering techniques on the efficacy of pixel-based encoders. These studies primarily utilize bidirectional encoders and process text as grayscale images.
@@ 2.1 Pixel Representations for Text
Models PIXEL PIXAR PixelGPT (Ours)
Image format Grayscale (0-1) Binary (0/1) RGB (0-255)
Modeling Bidirectional Autoregressive Autoregressive
Training Objective Regression Classification Regression
Modeling Space Continuous Discrete Continuous
Loss function Mean Squared Error Binary Cross Entropy Mean Squared Error
Table 1: Detailed comparison of pixel-based baselines.
For fair comparison, we summarize the comparison of our PixelGPT with pixel-based baselines, including PIXEL Rust et al. (2023), PIXAR Tai et al. (2024), in Table 1. It is worth noting that our work is different from PIXAR, which uses different training objectives and data rendering approaches from PIXEL and ours. Instead, our model can be seen as an autoregressive version of PIXEL.
2.2 Autoregressive Pre-training on Pixels
Existing methods in pixel-based autoregressive pre-training divide into vector quantization techniques—transforming continuous images into discrete tokens—and direct pixel prediction. These approaches include VQ-VAE Van Den Oord et al. (2017) and VQGAN Esser et al. (2021) followed by next token prediction Chen et al. (2020); Ramesh et al. (2021), and prefix language modeling that predicts future visual patches from bidirectional pixel contexts El-Nouby et al. (2024).
@@ 3.2 Input Representation
Image Input
Inspired by the Vision Transformer (ViT; Dosovitskiy et al., 2020), our method tailors the image patch processing paradigm to the sequential processing needs of autoregressive transformer decoders handling visual text imagery, as shown in Figure LABEL:fig:pixelgpt. This process commences by rendering the textual input as RGB images $x\in\mathbb{R}^{H\times W\times C}$ as aforementioned in §3.1, subsequently partitioning these into uniform patches $x_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$ illustrated as Figure 8, where $(H,W)$ defines the original image’s resolution, $(P,P)$ specifies each patch’s resolution with $P=H$, and $N=W/P$ denotes the total number of patches. The patches are then flattened, mapped to a $D$-dimensional space through a learnable linear projection, and finally fed into the transformer’s sequential processing stream. Unlike ViT, which caters to two-dimensional inputs, our model processes these patches in the sequence order in which the text appears, emulating the linear progression of reading. This patch-based segmentation aligns with the sequential nature of language, enabling our model to predictively learn from the visual data.
Text Input
We leverage the same tokenizer as Llama 2, segmenting input text into discrete tokens with a total vocabulary size of 32k. These tokens are then transformed into dense vector representations through an embedding lookup table.
@@ 3.3 Pre-training Objectives
Next Patch Prediction
Given a sequence of $N$ visual patches $x_{p}=(x^{1}_{p},x^{2}_{p},\cdots,x^{N}_{p})$ where each visual patch $x^{t}_{p}$ is a flattened patch embedding. We decompose the image patch sequence into the production of $N$ conditional probabilities:
$p(x^{1}_{p},x^{2}_{p},\cdots,x^{N}_{p})=\prod_{t=1}^{N}p(x^{t}_{p}|x^{1}_{p},x^{2}_{p},\cdots,x^{t-1}_{p})$
@@
Next Token Prediction
For text inputs, we utilize a conventional next token prediction objective, optimizing a cross-entropy loss that evaluates the fidelity of predicted token sequences generated via teacher-forcing against the ground truth tokens.
@@ 3.5 Pre-training Details
Model Architecture
Our architecture, illustrated in Figure LABEL:fig:arch, is built upon a stack of $N=24$ standard transformer decoder Vaswani et al. (2017), following Llama 2 Touvron et al. (2023b). We incorporate RMSNorm for pre-normalization Zhang and Sennrich (2019), SwiGLU activation functions Shazeer (2020); Chai et al. (2020), rotary position embeddings Su et al. (2024), and grouped query attention Ainslie et al. (2023). Comprehensive specifications and additional implementation details of our architecture are in Appendix §B.
Data
For visual image data, we use rendered the corpus of peS2o, English Wikipedia and C4 datasets for pre-training; while for text data, we adopt peS2o, English Wikipedia, C4, Common Crawl, and The Stack v1. We refer the readers to Appendix §C for details.
@@ 4.1 Experimental Setup
Baselines
For a thorough evaluation, we benchmark against models specialized in textual and visual representations. In the textual category, BERT and GPT-2 Radford et al. (2019) are chosen. For pixel-based models, we contrast our approach with DONUT Kim et al. (2022), CLIPPO Tschannen et al. (2023), and PIXEL Rust et al. (2023), which are trained on pixel-based representation. Detailed discussions are provided in Appendix §F.
@@ 4.2 Results
Autoregressive Pixel-based Pre-training Rivals PIXEL.
Our empirical investigation, detailed in Table 2, scrutinizes the feasibility of pure pixel-based autoregressive pre-training on RGB images of visual texts. The proposed PixelGPT model, training solely on rich raw visual inputs (24-bit RGB images), demonstrates not merely a competitive edge but, in several tasks, surpasses the performance of models pre-trained on text alone. Specifically, PixelGPT exhibits remarkable superiority on GLUE benchmarks – evidenced by its marked performance increases on the STS-B (+5.4), MRPC (+13.1), RTE (+11.9), and WNLI (+4.3) assessments compared to GPT-2. This demonstrates the viability of pixel-based pre-training in capturing complex linguistic constructs.
When compared to PIXEL, which leverages a bidirectional encoder architecture, PixelGPT exhibits enhanced performance in QQP (+1.5), RTE (+3.4), and WNLI (+5.4). These results collectively affirm the hypothesis that autoregressive pre-training on raw visual images is feasible for language modeling. PixelGPT achieves the optimal performance among pixel-based approaches on GLUE, underscoring the transformative impact of integrating rich visual information into pre-training. Refer to §G.5 for detailed discussion.
As shown in Figures 3 and 4, PixelGPT demonstrates a scaling trend with increased training data compute, indicating a promising direction for data scaling. This suggests that with more extensive training, PixelGPT has the potential to outperform text-based models, such as GPT-2 and BERT. Due to computational constraints, we will explore this in future work.
@@
Impact of Autoregressive Pixel Pre-training on Multilingual Tasks.
Synergistic Effects of Multimodal Pre-training.
@@
This comparative analysis, spanning both GLUE and XNLI datasets (the latter within the translate-train-all settings), is shown in Tables 4 and 5. A pivotal finding is that the incorporation of dual-modality data during pre-training markedly enhances average performance across language understanding tasks: DualGPT (76.9) surpasses both TextGPT (76.3) and MonoGPT (75.4). This suggests that potential conflicts arising from unimodal training can be significantly alleviated through a multimodal pre-training approach. This inference is corroborated by XNLI outcomes, wherein the addition of pixel-text paired data improved the model’s multilingual interpretative proficiency.
@@ 4.3 Analysis
Scaling Training Tokens vs. GLUE Performance
In Figure 3, we delineate the correlation between the scale of training data and the ensuing performance on the GLUE benchmark. Our analysis encompasses a spectrum of total training tokens/patches from 10 billion (B) to 240B, juxtaposing the trajectories of TextGPT, PixelGPT, MonoGPT, and DualGPT, with BERT and PIXEL serving as benchmarks. The MonoGPT and DualGPT models are evaluated under two different input modalities: text and pixel. From our findings, two primary insights emerge: (1) Pixel-based autoregressive pretraining models exhibit an increased data demand. With minimal training (e.g., at 10B), pixel-based models initiate at a lower performance threshold in pixel modality (all under 55%), compared to their text modality counterparts, which approximate a performance level of 70%. Nevertheless, with the increase of training data, a critical volume threshold catalyzes a substantial rise in performance for PixelGPT, MonoGPT, and DualGPT in pixel modality. This trajectory reveals a progressive convergence of PixelGPT towards the text-based baseline, culminating in its overtaking of PIXEL at around 200B tokens/patches and nearing TextGPT with a less than 5-point performance differential, while still on an upward trend. (2) The integration of paired dual-modality data during pretraining appears to confer significant benefits on multimodal learning, particularly for pixel-based input. When matched for training data volume, DualGPT consistently eclipses MonoGPT across comparable benchmarks, with the former maintaining a pronounced lead in pixel modality. This trend underscores the value of incorporating paired text-image data in pretraining to enhance the efficacy of multimodal learning.
@@
Scaling Training Tokens vs. XNLI (Translate-Train-All) Performance
We further explored the progression of model performance in multilingual capability across varying volumes of pre-trained tokens/patches. This comparison, delineated in Figure 4, focused on the Translate-Train-All setting of the XNLI benchmark.
(1) Pixel-based autoregressive models display a heightened requirement for training data in multilingual tasks, corroborating the trend observed on the GLUE benchmark. Initially, there is a notable performance disparity between pixel and text modalities, with pixel-based models lagging behind when training on a lesser volume of tokens/patches. However, this gap diminishes substantially with the increase in training volume. Remarkably, upon reaching the 200B, PixelGPT not only surpasses PIXEL but also matches the performance of BERT, indicating a continued potential for further enhancement in its multilingual proficiency with additional training data.
(2) The injection of dual-modality data at the early stages of training appears to be particularly beneficial for models learning from pixel data. When comparing DualGPT and MonoGPT under the pixel modality, DualGPT demonstrates a notable performance advantage at the outset of training (55% vs. 45.8% at the 10B token/patch mark). Although this edge tapers as the training volume expands, it suggests that early-stage multimodal alignment aids the pixel-based models in leveraging the textual data for enhanced multilingual understanding.
(3) Our text-based pre-training approach, TextGPT, demonstrates superior results over BERT. This is evident when training reaches approximately 100B tokens, where TextGPT outperforms BERT. This improvement may be attributed, in part, to our byte-level BPE tokenization as utilized in Llama 2, which effectively deconstructs unseen languages into their constituent raw bytes—a capability not afforded by BERT. Additionally, the enrichment of our text pre-training corpus from diverse sources contributes to this. For a detailed breakdown of the text pre-training data, we refer readers to Appendix §C.2.
A Large Batch Size Improves Stable Training
We observe a distinct preference for larger batch sizes when fine-tuning pixel-based modalities across certain datasets. As in Figure 5, we evaluate how different batch sizes—64, 128, 256, and 512—affect model performance on selected GLUE benchmark tasks, namely QQP, CoLA, and STS-B. A clear trend emerges from the data: increasing the batch size correlates with improved model performance. Our analysis suggests that pixel modality fine-tuning exhibits greater variance than text modality and benefits from the use of larger batch sizes. This appears to mitigate the variability inherent in different training batches, thus enhancing training stability. It prevents premature convergence to suboptimal local minima and fosters higher model accuracy.